OpenStudy (haseeb96):

There are 3 capacitors which draw current of 1A,4A and 2A respectively when connected across the terminals of battery.if they are connected in series with the same battery then the current will be drawn ? a. 2/7 A b.3/7 A c.4/7 A d.5/7 A e.9.7 A

OpenStudy (haseeb96):

@ParthKohli @robtobey @bohotness @butterflydreamer

OpenStudy (haseeb96):

@perl @AsiaJenell

OpenStudy (perl):

The total capacitance of capacitors in series is equal to the reciprocal of the sum of the reciprocals of their individual capacitances:

OpenStudy (haseeb96):

yes

OpenStudy (perl):

$$ \Large \frac{1}{C_\mathrm{total}} = \frac{1}{C_1} + \frac{1}{C_2} + \frac{1}{C_3} $$

OpenStudy (haseeb96):

but how can i calculate the total current ?

OpenStudy (perl):

you can plug in the individual capacitances in amps

OpenStudy (haseeb96):

how? can u give me an example?

OpenStudy (triciaal):